# RoboEXP: Action-Conditioned Scene Graph via Interactive Exploration for Robotic Manipulation

### v1

@@ Abstract
Robots need to explore their surroundings to adapt to and tackle tasks in unknown environments. Prior work has proposed building scene graphs of the environment but typically assumes that the environment is static, omitting regions that require active interactions. This severely limits their ability to handle more complex tasks in household and office environments: before setting up a table, robots must explore drawers and cabinets to locate all utensils and condiments. In this work, we introduce the novel task of interactive scene exploration, wherein robots autonomously explore environments and produce an action-conditioned scene graph (ACSG) that captures the structure of the underlying environment. The ACSG accounts for both low-level information, such as geometry and semantics, and high-level information, such as the action-conditioned relationships between different entities in the scene. To this end, we present the Robotic Exploration (RoboEXP) system, which incorporates the Large Multimodal Model (LMM) and an explicit memory design to enhance our system’s capabilities. The robot reasons about what and how to explore an object, accumulating new information through the interaction process and incrementally constructing the ACSG. We apply our system across various real-world settings in a zero-shot manner, demonstrating its effectiveness in exploring and modeling environments it has never seen before. Leveraging the constructed ACSG, we illustrate the effectiveness and efficiency of our RoboEXP system in facilitating a wide range of real-world manipulation tasks involving rigid, articulated objects, nested objects like Matryoshka dolls, and deformable objects like cloth. Project Page: https://jianghanxiao.github.io/roboexp-web/
{strip}
@@ I Introduction
Imagine a future household robot designed to prepare breakfast. This robot must efficiently perform various tasks such as conducting inventory checks in cabinets, fetching food from the fridge, gathering utensils from drawers, and spotting leftovers under food covers. Key to its success is the ability to interact with and explore the environment, especially to find items that aren’t immediately visible. Equipping it with such capabilities is crucial for the robot to effectively complete its everyday tasks.
Robot exploration and active perception have long been challenging areas in robotics [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Various techniques have been proposed, including information-theoretic approaches, curiosity-driven exploration, frontier-based methods, and imitation learning [17, 18, 19, 20, 21, 22, 1, 23, 14, 13, 24, 15, 25]. Nevertheless, previous research has primarily focused on exploring static environments by merely changing viewpoints in a navigation setting or has been limited to interactions with a small set of object categories, such as drawers, or a closed set of simple actions like pushing [26].
@@
In this work, we investigate the interactive scene exploration task, where the goal is to efficiently identify all objects, including those that are directly observable and those that can only be discovered through interaction between the robot and the environment (see Fig. 1). Towards this goal, we present a novel scene representation called action-conditioned 3D scene graph (ACSG). Unlike conventional 3D scene graphs that focus on encoding static relations, ACSG encodes both spatial relationships and logical associations indicative of action effects (e.g., opening a fridge will reveal an apple inside). We then show that interactive scene exploration can be formulated as a problem of action-conditioned 3D scene graph construction and traversal.
Tackling interactive scene exploration poses challenges: how can we reason about which objects need to be explored, choose the right action to interact with them, and maintain knowledge about our exploration findings? With these challenges in mind, we propose a novel, real-world robotic exploration framework, the RoboEXP system. At the core of our system is a large foundational model-powered instantiation of action-conditioned 3D scene graph. Specifically, our framework consists of four modules: perception, memory, decision-making, and action, as shown in Fig. 3. To address the challenge of perceiving what is present in the scene, our perception module utilizes Grounding-DINO ([27]), Segment Anything in High Quality (SAM-HQ) [28, 29], and CLIP [30] to detect objects or parts and extract their language-embedded semantic features. Our decision-making module employs the rich commonsense knowledge contained in large multimodal models, such as GPT-4V [31, 32], to assist in selecting which objects to explore and what actions to take, and in validating their plausibility. Once the decision-making module has chosen a skill, our action module is then activated to follow the plans formulated by the prior modules. During the entire physical interaction process, our memory model—which maintains the action-conditioned scene graph—will be continuously updated to preserve the scene’s knowledge for future exploration and exploitation. Despite its strong capacity, our hardware system is simple—it requires only a single RGB-D wrist camera as sensor input and uses a single robot arm for actions.
@@
RoboEXP can handle diverse exploration tasks in a zero-shot manner, constructing complex action-conditioned 3D scene graph in various scenarios, including those involving obstructing objects and requiring multi-step reasoning (Fig. 2). We evaluate our system across various settings, spanning simple, single-object scenarios to complex environments, demonstrating its adaptability and robustness. The system also effectively manages different human interventions. Moreover, we show that our reconstructed action-conditioned 3D scene graph demonstrates strong capacity in performing multiple complex downstream tasks. Action-conditioned 3D scene graph advances LLM/LMM-guided robotic manipulation and decision-making research [33, 34], extending their operation domain from environments with known or observable objects to complicated environments with unknown or unobserved ones. To our knowledge, this is the first of its kind.
Our contributions are as follows: i) we propose action-conditioned 3D scene graph and introduce the interactive scene exploration task to address the challenging interaction aspect of exploration; ii) we develop the RoboEXP system, capable of exploring complicated environments with unseen objects in a wide range of settings; iii) through extensive experiments, we demonstrate our system’s ability to construct complex and complete action-conditioned 3D scene graph, demonstrating significant potential for various manipulation tasks. Our experiments involve rigid and articulated objects, nested objects like Matryoshka dolls, and deformable objects like cloth, showcasing the system’s generalization ability across objects, scene configurations, and downstream tasks.
@@
Figure 2: Action-Conditioned 3D Scene Graph from Interactive Scene Exploration. To illustrate the construction process of our ACSG in the interactive scene exploration, we depict a scenario wherein a robot arm explores a tabletop scene containing two cabinets and a condiment obstructing the left door. (a) The robot arm actively interacts with the scene, completing the interactive scene exploration process. (b) We showcase the corresponding low-level memory in our ACSG, which represents the geometry and semantic information of the scene. The small graph within each visualization represents a segment of the final scene graph. (c) We present the high-level memory of our action-conditioned scene graph. The graph reveals that picking up the condiment serves as a precondition for opening the door, and opening the bottom drawer allows the observation of the concealed tape and banana.
II Related Works
Scene graphs [35, 36] represent objects and their relations [37, 38, 39] in a scene via a graph structure. Previous studies generate scene graphs from images [40, 36] or 3D scenes [41] with hierarchical and semantic information, and further with the assistance of large language models (LLMs) [42]. They leverage scene graphs for image captioning [43, 44], image retrieval and generation [35, 45], visual-language tasks [37, 46], navigation [47, 48] and task planning [49, 50]. While previous works model scene graphs in static 2D or 3D scenes, we generate action-conditioned scene graphs that integrate actions as core elements, depicting interactive relationships between objects and actions. This action-centric approach opens avenues for physical exploration and diverse downstream robotics tasks.
@@ II Related Works
Neuro-symbolic representations integrates neural networks’ perceptual abilities with the symbolic reasoning for robots in complex and dynamic environments. Prior works explored understanding scenes and describing robotic skills in symbolic texts to interpret demonstrations [51, 52], ground abstract actions for robotic primitives [53] and generate action plans [54, 55, 56, 57]. Our proposed framework also constructs symbolic representations of the environment, but in the form of action-conditioned scene graphs for robotic manipulation.
Robotic exploration aims to autonomously navigate, interact with, and gather information from environments it has never encountered before. It is applicable in search and rescue [1, 2, 58, 59, 60, 61, 62, 63, 64], planetary exploration [3, 4, 65, 66], object goal navigation [5, 6, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84], and mobile manipulation [7, 8, 85, 86, 87, 88]. The primary guiding principle behind robotic exploration is to reduce the uncertainty of the environment [89, 17, 58, 18, 19, 90], making uncertainty quantification key for robotic exploration tasks. Curiosity-driven exploration has recently emerged as a promising approach, showing effective results in various contexts  [15, 20, 21, 91]. Most past works have focused on exploration in the context of mobility [92, 1, 2, 58, 59, 60, 61, 62, 5, 6, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 7, 8, 85, 86, 87, 88], with the primary goal of modeling and understanding the static environment to complete specific tasks. Recently, exploration has also been studied in the context of manipulation [23, 93, 94, 16, 95], aiming to better understand the scene by changing the state of the environment. Our work introduces a new active exploration strategy for manipulation, uniquely defining a novel scene graph-guided objective to guide the exploration process.
@@
Active perception aims to select specific actions for an agent to improve its ability to perceive and understand the environment [9, 10]. Unlike passive perception, actions offer more flexibility, such as control over better viewpoints [11, 12, 13], sensor configurations [96, 14], or adjustments to environmental configurations [97]. It can also reveal certain scene properties that cannot be perceived in a passive manner, such as dynamic parameters [15, 24] or articulation [98, 16, 99]. Previous studies have explored active perception in 3D reconstruction [100, 101], object recognition [102, 103, 104], camera localization [105], and robotic manipulation [106, 107]. Our work falls into the category of actively exploring the environment to reveal what’s inside or underneath objects. Differing from most previous active perception efforts, which are driven by handcrafted rules [108], information gain [22, 109], or reinforcement learning [15, 25], our approach to active perception is guided by grounding the rich commonsense knowledge encoded in a large language model into an explicit scene graph representation.
@@
Language models for robotics. Large language models (LLMs) [110, 111, 112] and large multimodality models (LMMs) [31, 32] are bringing overwhelming influence into the robotics field, for their strong capacity in common-sense knowledge and long-horizon reasoning. Previous studies have harnessed the common-sense knowledge of such large models to generate action candidates [113] and action sequences for task planning [114, 112, 115, 84], and generate code for robotic control and manipulation [116, 33, 117]. More recently, VILA [34] utilized GPT-4V [31, 32] for vision-language planning. In our RoboEXP system, we leverage GPT-4V for decision-making in two crucial roles. First, as the action proposer, it ensures both effectiveness and efficiency in proposing appropriate strategies to expand potential nodes in our action-conditioned 3D scene graph. Second, as the action verifier, it ensures the plausibility and smoothness of actions and operations in our system. Moreover, instead of memorizing everything using large models in a brute force way, our system employs explicit memory to enhance the decision-making process.
Figure 3: Overview of Our RoboEXP System. We present a comprehensive overview of our RoboEXP system, comprised of four modules. (a) Our perception module takes RGBD images as input and produces the corresponding 2D bounding boxes, masks, object labels, and associated semantic features as output. (b) The memory module seamlessly integrates 2D information into the 3D space, achieving more consistent 3D instance segmentation. Additionally, it constructs the high-level graph of our ACSG through the merging of instances. (c) Our decision-making module serves dual roles as a proposer and verifier. The proposer suggests various actions, such as opening doors and drawers, while the verifier assesses the feasibility of each action, considering factors like obstruction. (d) The action module executes the proposed actions, enabling the robot arm to interact effectively with the environment.
III Problem Statement
We unfold this section with an introduction of action-conditioned 3D scene graph, a novel scene representation illustrating interactive object relationships (Sec. III-A). We then formulate interactive scene exploration as an action-conditioned 3D scene graph construction and traversal problem (Sec. III-B).
III-A Action-Conditioned 3D Scene Graph
@@
An action-conditioned 3D scene graph (ACSG) is an actionable, spatial-topological representation that models objects and their interactive and spatial relations in a scene. Formally, ACSG is a directed acyclic graph $\mathbf{G}=(\mathbf{V},\mathbf{E})$ where each node represents either an object (e.g., a door) or an action (e.g., open), and edges $\mathbf{E}$ represent their interaction relations. The object node $\mathbf{o}_{i}=(\mathbf{s}_{i},\mathbf{p}_{i})\in\mathbf{V}$ encodes the semantics and geometry of each object (e.g., the semantic embedding of a fridge $\mathbf{s}_{i}$, and its shape in the form of a point cloud $\mathbf{p}_{i}$), whereas the action node $\mathbf{a}_{k}=(a_{k},\mathbf{T}_{k})\in\mathbf{V}$ encodes high-level action type $a_{k}$ and low-level primitives $\mathbf{T}_{k}$ to perform the actions. Between the nodes are edges encoding their relations, which we categorize into four types: 1) between objects $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{o}}$ (e.g., the door handle belongs to the fridge), 2) from objects to actions $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{a}}$ (e.g., toy can be picked up), 3) from action to objects $\mathbf{e}_{\mathbf{a}\rightarrow\mathbf{o}}$ (e.g., a banana can be reached if we open the cabinet), or 4) from one action to another $\mathbf{e}_{\mathbf{a}\rightarrow\mathbf{a}}$ (e.g., the cabinet can be opened only if we move away the condiment). Our action-conditioned 3D scene graph greatly enhances existing 3D scene graphs, as it explicitly models the action-conditioned relations between objects. Fig. 2 depicts a complete action-conditioned 3D scene graph of a tabletop scene.
@@
One advantage of our interaction-aware scene graph lies in its simplicity for retrieving and taking actions on an object. Regardless of how complicated the scene is, given our scene graph and a target object, an agent merely needs to sequentially execute all the actions on the paths from the root to the object node in a topological order to retrieve the object. For example, in Fig. 2, to reach the tape inside a cabinet whose door is blocked by a condiment, according to the graph, one simply needs to: 1) pick up the condiment on the table that blocks the cabinet door, and 2) open the cabinet through the door handle.
III-B Interactive Exploration

### v2

@@ Abstract
We introduce the novel task of interactive scene exploration, wherein robots autonomously explore environments and produce an action-conditioned scene graph (ACSG) that captures the structure of the underlying environment. The ACSG accounts for both low-level information (geometry and semantics) and high-level information (action-conditioned relationships between different entities) in the scene. To this end, we present the Robotic Exploration (RoboEXP) system, which incorporates the Large Multimodal Model (LMM) and an explicit memory design to enhance our system’s capabilities. The robot reasons about what and how to explore an object, accumulating new information through the interaction process and incrementally constructing the ACSG. Leveraging the constructed ACSG, we illustrate the effectiveness and efficiency of our RoboEXP system in facilitating a wide range of real-world manipulation tasks involving rigid, articulated objects, nested objects, and deformable objects. Project Page: https://jianghanxiao.github.io/roboexp-web/
Keywords: Action-Conditioned Scene Graph, Foundation Models for Robotics, Scene Exploration, Robotic Manipulation
1 Introduction
@@
Imagine a household robot designed to prepare breakfast. This robot must perform various tasks such as conducting inventory checks in cabinets, fetching food from the fridge, gathering utensils from drawers, and spotting leftovers under food covers. Key to its success is the ability to interact with and explore the environment, especially to find items that aren’t immediately visible. Equipping it with such capabilities is crucial for the robot to effectively complete its everyday tasks. Robot exploration and active perception have long been challenging areas in robotics. Various techniques have been proposed, including information-theoretic approaches [1, 2, 3, 4, 5, 6, 7], frontier-based methods [8, 9, 10], imitation learning [11, 12] and reinforcement learning [13, 14, 15, 16, 17]. Nevertheless, previous research has primarily focused on exploring static environments by merely changing viewpoints in a navigation setting.
@@
In this work, we investigate the interactive scene exploration task, where the goal is to efficiently identify all objects, including those that are directly observable and those that can only be discovered through interaction between the robot and the environment (Fig. 1). Towards this goal, we present a novel scene representation called action-conditioned 3D scene graph (ACSG). Unlike conventional 3D scene graphs that focus on encoding static relations, ACSG encodes both spatial relationships and logical associations indicative of action effects (e.g., opening a fridge will reveal an apple inside). We then show that our task can be formulated as a problem of ACSG construction and traversal.
To tackle the interactive scene exploration task, we propose a novel, real-world robotic exploration framework, the RoboEXP system. At the core of our system is a large foundational model-powered instantiation of action-conditioned 3D scene graph. Specifically, our framework consists of four modules: perception, memory, decision-making, and action (Fig. 3).
@@
RoboEXP can handle diverse exploration tasks in a zero-shot manner, constructing complex action-conditioned 3D scene graph in various scenarios, including those involving obstructing objects and requiring multi-step reasoning (Fig. 2). We evaluate our system across various settings , demonstrating its adaptability and robustness. The system also effectively manages different human interventions. Moreover, we show that our reconstructed action-conditioned 3D scene graph demonstrates strong capacity in performing multiple complex downstream tasks. Action-conditioned 3D scene graph advances LLM/LMM-guided robotic manipulation and decision-making research [18, 19], extending their operation domain from environments with known or observable objects to complicated environments with unknown or unobserved ones. To our knowledge, this is the first of its kind.
@@
Figure 2: Action-Conditioned 3D Scene Graph from Interactive Scene Exploration. We depict a scenario wherein a robot arm explores a tabletop scene containing two cabinets and a condiment obstructing the left door. (a) The robot arm actively interacts with the scene, completing the interactive scene exploration process. (b) We showcase the corresponding low-level memory in our ACSG. The small graph on the bottom-left of each visualization represents a segment of the final scene graph. (c) We present the high-level memory of our ACSG. The graph reveals that picking up the condiment serves as a precondition for opening the door, and opening the bottom drawer allows the observation of the concealed banana.
2 Related Works
@@
Scene Graphs [20, 21] represent objects and their relations [22, 23, 24] in a scene via a graph structure. Previous studies generate scene graphs from images [25, 21, 26] or 3D scenes [27], and further with the assistance of large language models (LLMs) [28, 29]. While previous works model scene graphs in static 2D or 3D scenes, we generate action-conditioned scene graphs that integrate actions as core elements, depicting interactive relationships between objects and actions. Our work is also related to Neuro-Symbolic Representations that integrate neural networks with symbolic reasoning. Prior works explored understanding scenes and describing robotic skills in symbolic texts to interpret demonstrations [30, 31], ground abstract actions for robotic primitives [32] and generate action plans [33, 34, 35, 36, 37, 38]. Our proposed framework also constructs symbolic representations of the environment, but in the form of action-conditioned scene graphs for robotic manipulation.
@@
Robotic Exploration aims to autonomously navigate [9, 39, 40, 41, 42, 43, 11, 44, 45], interact with [12, 46, 46, 47, 48, 49, 50], and gather information [51, 52, 53] from environments it has never encountered before. The primary guiding principle behind robotic exploration is to reduce the uncertainty of the environment [54, 5, 39, 6, 7, 55], making uncertainty quantification key for robotic exploration tasks. Curiosity-driven exploration has recently emerged as a promising approach, showing effective results in various contexts  [16, 56, 57, 58]. Recently, exploration has also been studied in the context of manipulation [59, 60, 61, 62, 63, 64, 65, 66], aiming to better understand the scene by changing the state of the environment. Our work introduces a new active exploration strategy for manipulation, uniquely defining a novel scene graph-guided objective to guide the exploration process. Our work is also related to Active Perception, which actively selects actions for an agent to help it perceive and understand the environment [1, 67]. Unlike passive perception, actions offer more flexibility, such as control over viewpoints [2, 3, 4, 68], sensor configurations [69, 15], or adjustments to environmental configurations [70]. It can also reveal certain scene properties that cannot be perceived in a passive manner, such as dynamic parameters [16, 71] or articulation [72, 62, 73]. Our work falls into the category of actively exploring the environment to reveal what’s inside or underneath objects. Differing from most previous active perception efforts, which are driven by handcrafted rules [74], information gain [75, 76], or reinforcement learning [16, 77], our approach is guided by grounding the commonsense knowledge encoded in a LLM/LMM into an explicit scene graph representation.
@@
Language Models for Robotics. Large language models (LLMs) [78, 79, 80] and large multimodality models (LMMs) [81, 82] are bringing overwhelming influence into the robotics field, for their strong capacity in common-sense knowledge and task-level planning. Previous studies have harnessed the common-sense knowledge of such large models to generate action candidates [83, 84, 85] and action sequences for task planning [86, 80, 87, 88], and generate code for robotic control and manipulation [89, 18, 90]. More recently, VILA [19] utilized GPT-4V [81, 82] for vision-language planning. In our RoboEXP system, we leverage GPT-4V for decision-making in two crucial roles to select and verify actions. Moreover, instead of memorizing everything using large models in a brute force way, our system employs explicit memory to enhance the decision-making process.
3 Action-Conditioned 3D Scene Graph
@@
An action-conditioned 3D scene graph (ACSG) is an actionable, spatial-topological representation that models objects and their interactive and spatial relations in a scene. Formally, ACSG is a directed acyclic graph $\mathbf{G}=(\mathbf{V},\mathbf{E})$ where each node represents either an object or an action, and edges $\mathbf{E}$ represent their interaction relations. The object node $\mathbf{o}_{i}=(\mathbf{s}_{i},\mathbf{p}_{i})\in\mathbf{V}$ encodes the semantics $\mathbf{s}_{i}$ and geometry $\mathbf{p}_{i}$ of each object, whereas the action node $\mathbf{a}_{k}=(a_{k},\mathbf{T}_{k})\in\mathbf{V}$ encodes high-level action type $a_{k}$ and low-level primitives $\mathbf{T}_{k}$ to perform the actions. Between the nodes are edges encoding their relations, which we categorize into four types: 1) between objects $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{o}}$ (e.g., the door handle belongs to the fridge), 2) from objects to actions $\mathbf{e}_{\mathbf{o}\rightarrow\mathbf{a}}$ (e.g., toy can be picked up), 3) from action to objects $\mathbf{e}_{\mathbf{a}\rightarrow\mathbf{o}}$ (e.g., a banana can be reached if we open the cabinet), or 4) from one action to another $\mathbf{e}_{\mathbf{a}\rightarrow\mathbf{a}}$ (e.g., the cabinet can be opened only if we move away the condiment). Our action-conditioned 3D scene graph greatly enhances existing 3D scene graphs, as it explicitly models the action-conditioned relations between objects. Fig. 2 depicts a complete action-conditioned 3D scene graph of a tabletop scene.
@@
One advantage of our ACSG lies in its simplicity for retrieving and taking actions on an object. Regardless of how complicated the scene is, given our scene graph and a target object, an agent merely needs to sequentially execute all the actions on the paths from the root to the object node in a topological order to retrieve the object.